Using the the concept of rank of the matrix find whether the following system of equations

are consistent or inconsistent,

i.

4y + z = 0

12x - 5y - 3z = 34

-6x + 4z = 8

ii.

5x - 3y + z = 7

2x + 3y - z = 0

8x + 9y - 3z = 2:

iii.

-8x + 2z = 1

6y + 4z = 3

12x + 2y = 2


1
Expert's answer
2021-06-07T18:53:50-0400

a. "Ax=b" is inconsistent (i.e., no solution exists) if and only if "\text{rank}[A]<\text{rank}[A|b]."


b. "Ax=b" has an unique solution if and only if "\text{rank}[A]=\text{rank}[A|b]=n."


c. "Ax=b" has infinitely many solutions if and only if "\text{rank}[A]=\text{rank}[A|b]<n."


i.


"A=\begin{bmatrix}
 0 & 4 & 1 \\
 12 & -5 & -3 \\
 -6 & 0 & 4
\end{bmatrix}"

Swap rows 1and 2


"\begin{bmatrix}
 12 & -5 & -3 \\
 0 & 4 & 1 \\
 -6 & 0 & 4
\end{bmatrix}"

"R_3=R_3+(1\/2)R_1"

"\begin{bmatrix}
 12 & -5 & -3 \\
 0 & 4 & 1 \\
 0 & -5/2 & 5/2
\end{bmatrix}"

"R_3=R_3+(5\/8)R_2"


The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 3.


Since "\text{rank}[A]=\text{rank}[A|b]=3=n," then "Ax=b" has an unique solution.


ii.


"A=\begin{bmatrix}
 5 & -3 & 1 \\
 2 & 3 & -1 \\
 8 & 9 & -3
\end{bmatrix}"

"R_1=R_1\/5"

"\begin{bmatrix}
 1 & -3/5 & 1/5 \\
 2 & 3 & -1 \\
 8 & 9 & -3
\end{bmatrix}"

"R_2=R_2-2R_1"

"\begin{bmatrix}
 1 & -3/5 & 1/5 \\
 0 & 21/5 & -7/5 \\
 8 & 9 & -3
\end{bmatrix}"

"R_3=R_3-8R_1"

"\begin{bmatrix}
 1 & -3/5 & 1/5 \\
 0 & 21/5 & -7/5 \\
 0 & 69/5 & -23/5
\end{bmatrix}"

"R_2=(5\/21)R_2"

"\begin{bmatrix}
 1 & -3/5 & 1/5 \\
 0 & 1 & -1/3 \\
 0 & 69/5 & -23/5
\end{bmatrix}"

"R_1=R_1+(3\/5)R_2"

"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & -1/3 \\
 0 & 69/5 & -23/5
\end{bmatrix}"

"R_3=R_3-(69\/5)R_2"

"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & -1/3 \\
 0 & 0 & 0
\end{bmatrix}"

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 2.




"\begin{bmatrix}
 5 & -3 & 1 & & 7 \\
 2 & 3 & -1 & & 0 \\
 8 & 9 & -3 & & 2
\end{bmatrix}"

"R_1=R_1\/5"

"\begin{bmatrix}
 1 & -3/5 & 1/5 & & 7/5 \\
 2 & 3 & -1 & & 0 \\
 8 & 9 & -3 & & 2
\end{bmatrix}"

"R_2=R_2-2R_1"

"\begin{bmatrix}
 1 & -3/5 & 1/5 & & 7/5 \\
 0 & 21/5 & -7/5 & & -14/5 \\
 8 & 9 & -3 & & 2
\end{bmatrix}"

"R_3=R_3-8R_1"

"\begin{bmatrix}
 1 & -3/5 & 1/5 & & 7/5 \\
 0 & 21/5 & -7/5 & & -14/5 \\
 0 & 69/5 & -23/5 & & -46/5
\end{bmatrix}"

"R_2=(5\/21)R_2"

"\begin{bmatrix}
 1 & -3/5 & 1/5 & & 7/5 \\
 0 & 1 & -1/3 & & -2/3 \\
 0 & 69/5 & -23/5 & & -46/5
\end{bmatrix}"

"R_1=R_1+(3\/5)R_2"

"\begin{bmatrix}
 1 & 0 & 0 & & 1 \\
 0 & 1 & -1/3 & & -2/3 \\
 0 & 69/5 & -23/5 & & -46/5
\end{bmatrix}"

"R_3=R_3-(69\/5)R_2"

"\begin{bmatrix}
 1 & 0 & 0 & & 1 \\
 0 & 1 & -1/3 & & -2/3 \\
 0 & 0 & 0 & & 0
\end{bmatrix}"

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 2.

Since "\text{rank}[A]=\text{rank}[A|b]=2<3=n," then "Ax=b" has infinitely many solutions.


iii.



"A=\begin{bmatrix}
 -8 & 0 & 2 \\
 0 & 6 & 4 \\
 12 & 2 & 0
\end{bmatrix}"

"R_1=-R_1\/8"

"\begin{bmatrix}
 1 & 0 & -1/4 \\
 0 & 6 & 4 \\
 12 & 2 & 0
\end{bmatrix}"

"R_3=R_3-12R_1"

"\begin{bmatrix}
 1 & 0 & -1/4 \\
 0 & 6 & 4 \\
 0 & 2 & 3
\end{bmatrix}"

"R_2=R_2\/6"

"\begin{bmatrix}
 1 & 0 & -1/4 \\
 0 & 1 & 2/3 \\
 0 & 2 & 3
\end{bmatrix}"

"R_3=R_3-2R_2"

"\begin{bmatrix}
 1 & 0 & -1/4 \\
 0 & 1 & 2/3 \\
 0 & 0 & 5/3
\end{bmatrix}"

"R_3=(3\/5)R_3"

"\begin{bmatrix}
 1 & 0 & -1/4 \\
 0 & 1 & 2/3 \\
 0 & 0 & 1
\end{bmatrix}"

"R_1=R_1+(1\/4)R_3"

"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 2/3 \\
 0 & 0 & 1
\end{bmatrix}"

"R_2=R_2-(2\/3)R_3"

"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1
\end{bmatrix}"


The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 3.


Since "\text{rank}[A]=\text{rank}[A|b]=3=n," then "Ax=b" has an unique solution.